OpenStudy (anonymous):

Differentiate cos^2x - sin^2x

jigglypuff314 (jigglypuff314):

chain rule and know that derivative of (cos(x)) = -(sin(x)) and derivative of (sin(x)) = (cos(x)) derivative of (stuff)^2 = 2(stuff) * deriv(stuff)

jigglypuff314 (jigglypuff314):

you can think of cos^2x - sin^2x as (cos(x))^2 - (sin(x))^2 now use chain rule so 2(cos(x)) * deriv(cos) - 2(sinx) * deriv(sinx) follow me so far?

OpenStudy (anonymous):

Thanks, yeah kind of

jigglypuff314 (jigglypuff314):

then 2(cos(x)) * deriv(cos) - 2(sinx) * deriv(sinx) ^=(-sinx) ^=(cosx) so you get 2(cosx)(-sinx) - 2(sinx)(cosx)

jigglypuff314 (jigglypuff314):

2(cosx)(-sinx) - 2(sinx)(cosx) -> -2(sinx)(cosx) - 2(sinx)(cosx) and just like how -2x -2x = -4x -2(sinx)(cosx) - 2(sinx)(cosx) = -4(sinx)(cosx)

OpenStudy (anonymous):

Ahh okay thank you. That makes sense now

jigglypuff314 (jigglypuff314):

you're welcome! glad I could help! :D

terenzreignz (terenzreignz):

All's well, except -4(sin x)(cos x) could be further simplified... A lot could be revealed by just remembering that \[\Large \cos^2(x) -\sin^2(x) = \cos(2x)\]